OpenStudy (anonymous):

if you have a helment is 28.00 and you have a discount of 25% off then you must pay 6% sales tax find the total cost for the bike helment

OpenStudy (jdoe0001):

\(\bf {\color{red}{ a}}\%\ of \ b=\cfrac{{\color{red}{ a}}}{100}\times b\\ \quad \\ \quad \\ {\color{red}{ 25}}\%\ of \ 28=\cfrac{{\color{red}{ 25}}}{100}\times 28\\ \quad \\ \textit{so the helmet discount price is }28-\left(\cfrac{{\color{red}{ 25}}}{100}\times 28\right)\\ \quad \\ {\color{red}{ 6}}\%\ of \ \left[28-\left(\cfrac{{\color{red}{ 25}}}{100}\times 28\right)\right]=\cfrac{{\color{red}{ 6}}}{100}\times \left[28-\left(\cfrac{{\color{red}{ 25}}}{100}\times 28\right)\right]\\ \quad \\ \textit{so the total cost of the helmet is }\\ \quad \\28-\left(\cfrac{{\color{red}{ 25}}}{100}\times 28\right)+\left[\cfrac{{\color{red}{ 6}}}{100}\times \left[28-\left(\cfrac{{\color{red}{ 25}}}{100}\times 28\right)\right]\right]\)

OpenStudy (anonymous):

whats the awnser